# Prediction-Powered Adaptive Shrinkage Estimation

### v1

@@ D.5 Benchmarking in real-world datasets
Let $K$ be the number of experiment trials, $T_{j}$ be the total number of data points for problem $j$, i.e. $\{\dot{X}_{ij},\dot{Y}_{ij}\}_{j=1}^{T_{j}}$ represents the “raw data” we have, and $n_{j},N_{j}$ be the desired number of labeled/unlabeled data to simulate, usually calculated through a hyper-parameter splitting ratio (e.g. $N_{j}=\lfloor r\cdot T_{j}\rfloor,\>n_{j}=T_{j}-N_{j}$ for $r=0.8$ in our case).
@@
    Following evaluation methodology in existing PPI literature, e.g., [Angelopoulos et al., 2023], we first calculate the mean of all responses for each problem and treat it as the pseudo ground-truth, i.e., $\dot{\theta}_{j}:=\frac{1}{T_{j}}\sum_{i}\dot{Y}_{ij}$.
@@
    $\widehat{\text{MSE}}_{K}(\boldsymbol{\hat{\theta}})$ $:=\frac{1}{K}\sum_{k=1}^{K}\left(\frac{1}{m}\sum_{j=1}^{m}(\hat{\theta}_{j}^{k}-\dot{\theta}_{j})^{2}\right),$
    $\text{SE}_{K}(\boldsymbol{\hat{\theta}})$ $:=\frac{1}{\sqrt{K}}\sqrt{\frac{1}{K-1}\sum_{k=1}^{K}\left(\frac{1}{m}\sum_{j=1}^{m}(\hat{\theta}_{j}^{k}-\dot{\theta}_{j})^{2}-\widehat{\text{MSE}}_{K}(\boldsymbol{\hat{\theta}})\right)}.$
It should be noted that the standard error only accounts for uncertainty due to the random splits into labeled and unlabeled datasets.
D.6 Computational Resources
@@
D.7 Code Availability
The code for reproducing the experiments is available at

### v2

@@ E.4 Benchmarking in real-world datasets
Let $K$ be the number of experiment trials, $T_{j}$ be the total number of data points for problem $j$, i.e. $\{\dot{X}_{ij},\dot{Y}_{ij}\}_{i=1}^{T_{j}}$ represents the “raw data” we have, and $n_{j},N_{j}$ be the desired number of labeled/unlabeled data to simulate, usually calculated through a hyper-parameter splitting ratio (e.g. $N_{j}=\lfloor r\cdot T_{j}\rfloor,\>n_{j}=T_{j}-N_{j}$ for $r=0.8$ in our case).
@@
    Following evaluation methodology in existing PPI literature, e.g., (Angelopoulos et al., 2023), we first calculate the mean of all responses for each problem and treat it as the pseudo ground-truth, i.e., $\dot{\theta}_{j}:=\frac{1}{T_{j}}\sum_{i}\dot{Y}_{ij}$.
@@
    $\widehat{\text{MSE}}_{K}(\boldsymbol{\hat{\theta}})$ $:=\frac{1}{K}\sum_{k=1}^{K}\left(\frac{1}{m}\sum_{j=1}^{m}(\hat{\theta}_{j}^{k}-\dot{\theta}_{j})^{2}\right),\quad\text{SE}_{K}(\boldsymbol{\hat{\theta}}):=\frac{1}{\sqrt{K}}\sqrt{\frac{1}{K-1}\sum_{k=1}^{K}\left(\frac{1}{m}\sum_{j=1}^{m}(\hat{\theta}_{j}^{k}-\dot{\theta}_{j})^{2}-\widehat{\text{MSE}}_{K}(\boldsymbol{\hat{\theta}})\right)}.$
Note that the standard error only accounts for uncertainty due to the random splits into labeled and unlabeled datasets.
E.5 Additional Experiment with Varying Labeled/Total Data Ratio
In Table 3, we report our experiment results on different tasks and predictors, but fixing the ratio between labeled and total amounts of data $n_{j}/(n_{j}+N_{j})=0.2$ for all problems. To verify the broader applicability of our method, we repeat our experiments across a much wider range of ratios—from 1% to 40%—and report the results in Figure 6. For each ratio, we follow exactly the same data-splitting and benchmarking procedures specified in Section E.4.
Figure 6: Average MSEs for the three real-world datasets when the labeled/unlabeled split ratio varies from 1% to 40%. UniPT and UniPAS are the two newly added variants of PT and PAS estimators, respectively.
E.6 Computational Resources
@@
E.7 Code Availability